OpenStudy (anonymous):

4x=8y+4 5x-3y=5 Substitution method.

OpenStudy (kira_yamato):

4x=8y+4 => x=2y+1 ---(1) 5x-3y=5 --- (2) Sub (1) into (2): 5(2y+1)-3y = 5 10y-3y+5=5 7y=0 => y=0 Sub y=0 into (1): x=1

OpenStudy (anonymous):

Whew, I was worried I did something wrong with that one. Thank you.